the difference of two numbers is 36. one sixth of the larger number minus one ninth of the smaller number is 11. find the numbers
Answer by Kingsley(15)

X - y = 36---(1)
1/6X - 1/9Y = 11---(2)
X = 36+Y---(3)
1/6(36+Y)-1/9Y = 1
36+Y/6 - Y/9 = 1
108 + 3Y-2Y/18 = 1
108 + y = 18
y = 90
finding x,go to eqn(1)
x = 36+90
x = 126.
Thus,x=126 nd y = 90